19) (9 pts) Identify the relationship between each pair of structures shown A) same compound, B) constitutional isomers, C) conformational isomers, D) configurational isomers or E) enantiomers. CH B CH and Answer *To receive full credit, label each as R or S. b. Answer: Cl Br CH3 *To receive full credit, label each as R or S. c. Answer o (5 pts) Determine the formal charge on the indicated atoms. HzN-N a. : Br: b.
Answer #1

19.

Constitutional Isomers: Constitutional isomers possess different bond connectivities and have different physical and chemical properties.

The constitution of a compound defines the nature and sequence of bonding of the atoms. Isomers differing in constitution are termed constitutional isomers. Ethyl methyl ether and 1-propanol are constitutional isomers.

Stereoisomers: Stereoisomers are isomers with the same constitution, but differ in the arrangement of their atoms in space. They may have different physical and chemical properties.

Stereoisomers can be defined further into conformers and configurational isomers .

Conformational Isomers: Conformational isomers (conformers) are structures that differ only by rotation about a single bond. They cannot normally be separated and possess identical physical and chemical properties.

No bonds are broken when one conformation is converted into another.

Configurational Isomers: Configurational isomers differ in the arrangements of atoms in space, but excludes conformers. To change one configurational isomer into another requires breaking of bonds.

Configurational isomers can be conveniently divided into two classes: enantiomers and diastereoisomers (also called diastereomers).

Diastereoisomers: Diastereoisomers are configurational isomers that are not mirror images of one another.

Enantiomers: An enantiomer is a configurational isomer that is non-superposable on its mirror image.

20. Formal charge (FC):

In chemistry, a formal charge (FC) is the charge assigned to an atom in a molecule, assuming that electrons in all chemical bonds are shared equally between atoms, regardless of relative electronegativity.

Formal Charge (FC) = No. of valency electrons- (bond+dots)

(a) F.C of Na = 5-(2+4) = -1 (valency of nitrogen is 5)

F.C of Nb = 5- (4-0) = +1

F. C of Nc = 5-(3+2) = 0

(b) F.C of Ca = 4-(3+2) = -1 (valency of carbon is 4)

F.C of Brb = 7- (1+6) = 0 (valency of bromine is 7)